HARPS News Updates

On this page, we list the main changes that affected the instrument and this web site. More News and Information can be found in the Science information pages

  • Nov. 8, 2021: The main mirror of the telescope was successfully re-aluminized. The telescope is in a very good shape and is back to operations.
  • Nov. 8, 2021: the lowest hatch of the dome cannot be moved until repair. From today the limiting airmass for observations will be 1.7. Objects lower than this will suffer vignetting.
  • Jan 15, 2021: Personnel shuffle. As Florian Rodler took over the instrument CRIRES on Paranal, Gaspare LoCurto is back as instrument scientist for HARPS (and NIRPS).
  • Jan 4, 2021: the retarder plates rotator is currently out of operations. This means that the polarimetry mode cannot be offered at the moment.
  • Nov 26, 2020: due to the restricted-mode operations, we do not accept DDT proposals until further notice.
  • Nov 9, 2020: HARPS is back in operations. IMPORTANT:  You will have to account for an RV offset value (i.e. you add this as a free parameter to your RV fit) for data taken prior and after 2020-03-23. Nota bene: such an RV offset is always necessary when changes in the instrument (e.g. fiber exchange, warm-up) or data-reduction software were done, and your data set comprises of observations prior and after such an intervention. Such dates are 2015-06-02 (fiber exchange) and 2020-03-23 (warm-up of HARPS due to the COVID-19 pandemic).
  • Mar 22, 2020. Due to the COVID-19 pandemic, science operations at the observatory are stopped However the HARPS detectors will be kept cold.
  • Mar 11, 2020: the problem regarding lower flux in the calibration fiber at high airmasses has been fixed.
  • Mar 3, 2020: simultaneous FP and ThAr flux lower at high airmasses. We have noticed that the HARPS calibration fiber exhibits less flux when the telescope is very inclined (high airmasses). We are investigating the cause of the technical problem.

  • Nov 28, 2019: The NIRPS front-end hosting the adaptive optics of this new instrument has been installed above the HARPS telescope adapter. In this way simultaneous operations of HARPS+NIRPS will be possible. This change has no effect on the RV precision and has not introduced any RV offset.
  • Oct 12, 2019: P2 is online (https://www.eso.org/p2ls/). Please use P2 to create your OBs. (P2PP has been decommmissioned.)
    ==> If you wish to grant access to your OBs to an observer, please login into your ESO account via the ESO User Portal. There, click on 'Delegate Phase 2 tasks'. Select the run you wish to delegate, enter the email address of the ESO account associated with the observer and tick 'Phase 2'. Now the observer will be able to see, edit and execute the OBs of your run from the ESO account associated with her/him. Please note that delegating your run to another person / ESO account will not change the ID of your observing run.
  • IMPORTANT - Sep 27, 2019: P2PP goes P2. Please note that due to a software upgrade on La Silla on Oct 12, 2019, P2PP will cease to work. For all observing runs after Oct 12, 2019, please use the web-based La Silla P2 tool (https://www.eso.org/p2ls/) to prepare the observations. Your OBs are then automatically transferred to La Silla; no need to ftp them anymore.
  • Sep 27, 2019: Pipeline problems affecting data taken with the Fabry Perot in the time frame 02-Feb-2019 to 16-Apr-2019 - DONE. Due to a problem in the HARPS pipeline recipe the Fabry Perot calibration files after 01-Feb-2019 were not copied into the calibration database, and therefore the reference Fabry Perot spectrum was not updated. For this reason, all night-time science data taken with the Fabry Perot in the time frame 02-Feb-2019 to 16-Apr-2019 at 13:14 UT were reprocessed with the correct calibration files and are available in the archive.

  • November 15, 2018: The Laser Frequency Comb (LFC) is now running at HARPS. It is used for calibrations only and is not offered for night-time observations.
  • February 23, 2016: A Messenger article describing the May 2015 upgrade to the fiber-link is available in the December issue.
  • August 1, 2015: A letter has been sent to all HARPS Programme PI about the upgrade mentioned earlier and is also available here.
  • July 4, 2015: The HARPS polarimeter has now been realigned and the flux on both fibres is now within 30% (fibre B being less efficient).
  • June 3, 2015: After 12 years of operations a new fiber-link has been installed on HARPS. This includes the EGGS mode, and  the "sky fiber" of this mode is now available. Therefore starting from June 1st 2015 we expect a change in zero point in the determination of the radial velocities. We have collected the data to determine this offset as a function of stellar spectral type, and we are going to make the values available soon. We expect the effect of de-focussing and de-centering on the RV determination to be strongly reduced. The polarimetric mode of HARPS is not yet realigned with the new fibers and we measure a lower flux on fiber B by a factor 2 - 3. We intend to recover this as soon as possible. A full report will be made available to the community through these web pages and later through the ESO messenger.
  • October 21st, 2014: HARPS Pipeline Processed Data Available through Phase 3
  • April 6th, 2011: User Manual updated to include information on the use of the FP and a brief summary of the upgrade of the 3.6-m guiding system.
  • October 20th, 2010: New pipeline release: 3.5
  • March 2010 The commissioning data of the new spectro-polarimetric mode are released to the astronomical community. Link to this page to access them.
  • February 15th, 2010 A spectropolarimeter is now installed in HARPS. See the ESO announcement
  • October 17th, 2008: New pipeline release: 3.4
  • August 22nd, 2008: New flat field lamp (with filters) installed. More uniform spectrum across the HARPS wavelength range.
  • August 20th, 2008: The EGGS system runs in single fiber mode only. The reference fiber is damaged. However, the main science mode of EGGS is unaffected. Read more.
  • May 1st, 2007: HARPS reduced data can be requested from the HARPS Advanced Data Products archive.
  • December 08th, 2006: New pipeline release: 3.0.
  • May 15th, 2006: New, "slow" readout mode available.
  • May 15th, 2006: New release of the User manual for P78
  • March 1st, 2006: The high efficiency fibre link (EGGS) has been included in the call for proposal for period 78.
  • August 8, 2004: Secondary mirror unit changed; system efficiency increased by up to 40%.
  • May, 2004: Iodine self calibration method not available.
  • December, 2003: First HARPS planet candidate (HD330075b): Announcement, Paper
  • October 1, 2003: Instrument offered to the community.
  • February 11, 2003: FIRST LIGHT! ESO Press Release (06/03)

Press Releases

HARPS in the general press